solve the equation if possible:
2x +7*sqr(x) = 4

Answer by rothauserc(4718)

2x +7*sqr(x) = 4
:
7 * sqrt(x) = 4 - 2x
:
square both sides of =
:
49x = 16 -16x + 4x^2
:
4x^2 -65x +16 = 0
:
using quadratic formula, the discriminant is + or - 63
:
*****************************************
x = (-(-65) + 63) / (2 * 4) = 16
x = (-(-65) - 63) / (2 * 4) = 1/4 = 0.25
*****************************************
: